OpenStudy (anonymous):

What are the zeros of f(x) = x^3 + 6x^2 + 3x – 10 ?

OpenStudy (lgbasallote):

use rational roots theorem

OpenStudy (anonymous):

When I do that I get \[\pm1.\pm2,\pm5, and \pm10\] But there are only 3 zeros. What am I doing wrong?

OpenStudy (lgbasallote):

those are POSSIBLE roots...they're not automatically the zeros

OpenStudy (lgbasallote):

so you have to try them one by one...three of them are the roots to the equation

OpenStudy (lgbasallote):

do you want to know a shortcut?

OpenStudy (anonymous):

I would love to know a shortcut!

OpenStudy (lgbasallote):

pick a number from those numbers you listed

OpenStudy (anonymous):

\[\pm2 \]

OpenStudy (lgbasallote):

that means +2 and -2 <--those are two numbers...pick one of those

OpenStudy (anonymous):

Okay, +2.

OpenStudy (lgbasallote):

okay...now substitute +2 into EVERY x of x^3 + 6x^2 + 3x - 10 then tell me what the result is

OpenStudy (anonymous):

I get to 28.

OpenStudy (lgbasallote):

IF you get a result that is 0 then that is the ZERO..in this case you got 28 so it's not a zero

OpenStudy (anonymous):

Other way is to jus factorize. A factor of x+a implies -a is a zero.

OpenStudy (lgbasallote):

for example.. +1 (1)^3 + 6(1)^2 + 3(1) - 10 = 1 + 6(1) + 3 - 10 = 4 + 6 - 10 = 10 - 10 = 0 therefore x = 1 is a zero i gave you a bonus already...can you solve the other two zeroes now? ;)

OpenStudy (anonymous):

I have it from here, thank you. I've got 1, and -2 so far.

OpenStudy (lgbasallote):

welcome ^_^ and nice going

OpenStudy (lgbasallote):

actually...since it's a cubic function, once you get a zero you can just divide to get a quadratic equation then just factor